OpenStudy (anonymous):

square root of 512k square

OpenStudy (anonymous):

could you be more specific? your numbers are confusing

jhonyy9 (jhonyy9):

wann being sqrt(512k^2) =

OpenStudy (anonymous):

\[\sqrt{512*k ^{2}}=k*\sqrt{512}\]

OpenStudy (anonymous):

\[k*\sqrt{512}=k*\sqrt{2*256}=k*16*\sqrt{2}\]

jhonyy9 (jhonyy9):

nice